40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 1 08:50:48 2022, yuta, Summary, LSC, FPMI with REFL/AS55 trial 
    Reply  Thu Jul 14 00:10:08 2022, yuta, Summary, LSC, FPMI with REFL/AS55 trial continued AS55_I.pngAS55_Q.pngREFL55_I.pngREFL55_Q.png
       Reply  Fri Jul 15 19:13:22 2022, Paco, Summary, LSC, FPMI with REFL/AS55 demod phase adjust 
          Reply  Fri Jul 15 22:36:04 2022, rana, Summary, LSC, FPMI with REFL/AS55 demod phase adjust 
          Reply  Mon Jul 18 16:39:07 2022, Paco, Summary, LSC, FPMI locking procedure using REFL55 and AS55 DARM_07_18_2022_FMPI.pdfCARM_07_18_2022_FPMI.pdfMICH_07_18_2022_FPMI.pdffpmi_darm_nb_2022_07.pdf
             Reply  Mon Jul 18 21:41:42 2022, Anchal, Summary, LSC, FPMI locking procedure using REFL55 and AS55 
             Reply  Tue Aug 9 19:54:31 2022, yuta, Summary, LSC, FPMI locking tonight 
                Reply  Thu Aug 18 14:49:35 2022, Yehonathan, Summary, LSC, FPMI Sensitivity Sensitivity_Plot_1344133503.pdf
                   Reply  Thu Aug 18 18:10:49 2022, Koji, Summary, LSC, FPMI Sensitivity 
Message ID: 17089     Entry time: Thu Aug 18 14:49:35 2022     In reply to: 17069     Reply to this: 17091
Author: Yehonathan 
Type: Summary 
Category: LSC 
Subject: FPMI Sensitivity 

{Yuta, Yehonathan}

We wrote a notebook found on Git/40m/measurements/LSC/FPMI/NoiseBudget/FPMISensitivity.ipynb for calculating the MICH, DARM (currently XARM), CARM (currently YARM) sensitivities in the FPMI lock which can be run daily.

The IN and OUT channels of each DOFs are measured at a certain GPS time and calibrated using the optical gains and actuation calibration measured in the previous post.

Attachment shows the results.

It seems like the UGFs for MICH and DARM (currently XARM) match the ones that were estimated previously (100Hz for MICH, 120Hz for DARM) except for CARM for which the UGF was estimated to be 250Hz and here seems to be > 1kHz.

Indeed one can also see that the picks in the CARM plot don't match that well. Calculation shows that at 250Hz OUT channel is 6 times more than the IN channel. Calibrations for CARM should be checked.

MICH sensitivity using REFL55 at high frequencies is not much better than what was measured with AS55.

DARM sensitivity at 10Hz is a factor of a few better than the single arm lock sensitivity.

Now it is time to do the budgeting.

Attachment 1: Sensitivity_Plot_1344133503.pdf  285 kB  Uploaded Thu Aug 18 16:15:08 2022  | Hide | Hide all
Sensitivity_Plot_1344133503.pdf
ELOG V3.1.3-